# Learning to Steer Markovian Agents under Model Uncertainty

### v2

@@ Acknowledgement
The codes for all the experiments in this paper and the instructions for running can be found in the supplementary materials.
@@ Appendix E Missing Proofs for Existence when the True Model $f^{*}$ is Unknown
Consider the steering strategy $\psi({\bm{\pi}})=u_{\bm{\pi}}$. Given any $f\in\mathcal{F}$, and the trajectory sampled by $\psi$ and $f$, by Lem. E.2, w.p. $1-\frac{\delta}{|\mathcal{F}|}$, we have:
@@
By union bound, if $T_{0}=\lceil\frac{4}{\zeta}\log\frac{|\mathcal{F}|}{\delta}\rceil+1$, with probability at least $1-\delta$,
@@ G.5 A Summary of the Compute Resources by Experiments in this Paper
Appendix H Discussion on Limitations and Societal Impact
Discussion on Limitations
The main limitation of this paper is that we only focus on Markovian agents, whose dynamics $f$ only depends on their current policy profiles and the (modified) reward function. Although this model can already capture lots of learning dynamics, not all the theoretical results and methodology can be generalized to non-Markovian settings.
Besides, we can not guarantee how the optimal steering strategy regarding our Objective (1) performs in terms of the steering cost. We leave it to the future works.
Moreover, for the large model class setting, the strategy obtained by running proposed FETE framework may not have guarantees on the steering cost. It is also unclear if there are better algorithm designs for this setting.
Discussion on Societal Impact
Although this work are not directly applicable to real-world applications, but we believe it may have positive societal impact. The steering problem we study matches the requirement of many real-world scenarios (e.g. the social planner may incentivize technology companies by individual financial subsidies). We believe the theoretical and algorithmic contributions in this paper can provide useful insights for those applications.

### v3

@@ Acknowledgement
The code of all the experiments in this paper and the instructions for running can be found in https://github.com/jiaweihhuang/Steering_Markovian_Agents.
@@ Appendix C Missing Proofs in Section 3
Remark C.1.
Intuitively, Pareto optimality ensures the steering rewards predicted by $\psi^{*}$ are not necessarily large. For example, consider agents following NPG dynamics (Def. 4.1). Given any steering strategy $\psi$ and a constant $c>0$, consider another steering strategy $\psi^{\prime}(\cdot):=\psi(\cdot)+c$, i.e., by increasing the steering reward predicted by $\psi$ with $c$. Note that the learning dynamics under both $\psi$ and $\psi^{\prime}$ are the same, and therefore the final steering gaps are the same. Pareto optimality implies that $\psi^{\prime}$ will never be more preferable than $\psi$.
@@ Appendix E Missing Proofs for Existence when the True Model $f^{*}$ is Unknown
Consider the steering strategy $\psi({\bm{\pi}})=u_{\bm{\pi}}$. Given any $f\in\mathcal{F}$, and the trajectory sampled by $\psi$ and $f$, by Lem. E.2, w.p. $1-\frac{\delta}{|\mathcal{F}|}$, if $f_{\text{MLE}}\neq f$, we have:
@@
By union bound over all $f\in\mathcal{F}$, if $T_{0}=\lceil\frac{4}{\zeta}\log\frac{|\mathcal{F}|}{\delta}\rceil+1$, the following holds:
@@ G.5 A Summary of the Compute Resources by Experiments in this Paper
Appendix H Additional Discussion about Generalizing our Results
In this section, we discuss some extensions of the principle and algorithms in this paper to more general settings.
Non-Tabular Setting
When the game is non-tabular and its state and action spaces are infinite, the steering problem itself is fundamentally challenging without additional assumptions, since the policies are continuous distributions with infinite dimension and the learning dynamics can be arbitrarily complex. Nonetheless, when the agents’ policies and steering rewards are parameterized by finite variables, our methods and algorithms can still be generalized by treating the parameters as representatives.
As a concrete example, the Linear-Quadratic (LQ) game is a popular model involving countinuous state and action spaces (Jacobson,, 1973; Başar and Bernhard,, 2008; Zhang et al.,, 2019). In zero-sum LQ game, the game dynamics are characterized by a linear system:
$x_{t+1}=Ax_{t}+By_{t}+Cz_{t},$
with one-step reward function
$r^{1}(x_{t},y_{t},z_{t})=-r^{2}(x_{t},y_{t},z_{t})=x_{t}^{\top}Qx_{t}+y_{t}^{\top}R^{u}y_{t}-z_{t}^{\top}R^{v}z_{t}.$
Here $x_{t},x_{t+1}\in{\mathbb{R}}^{d}$ are the system states, $y_{t}\in{\mathbb{R}}^{m_{1}}$ and $z_{t}\in{\mathbb{R}}^{m_{2}}$ denote the actions of two agents.
Besides, the agents policies are parameterized by matrices $K_{t}\in{\mathbb{R}}^{m_{1}\times d},L_{t}\in{\mathbb{R}}^{m_{2}\times d}$, i.e.
$y_{t}=-K_{t}x_{t},\quad z_{t}=-L_{t}x_{t}.$
Following the quadratic form of the original reward, one may consider quadratic steering reward functions with parameters $\Theta_{t}:=(\Theta_{t}^{Q},\Theta_{t}^{u},\Theta_{t}^{v})$ and $\Xi_{t}:=(\Xi_{t}^{Q},\Xi_{t}^{u},\Xi_{t}^{v})$, such that, the steering reward for two agents at step $t$ is specified by:
$u^{1}_{t}(x_{t},y_{t},z_{t})=x_{t}^{\top}\Theta_{t}^{Q}x_{t}+y_{t}^{\top}\Theta_{t}^{u}y_{t}-z_{t}^{\top}\Theta_{t}^{v}z_{t},$
$u^{2}_{t}(x_{t},y_{t},z_{t})=x_{t}^{\top}\Xi_{t}^{Q}x_{t}+y_{t}^{\top}\Xi_{t}^{u}y_{t}-z_{t}^{\top}\Xi_{t}^{v}z_{t},$
and the reward after modification would be:
$r^{1}(x_{t},y_{t},z_{t})+u^{1}_{t}(x_{t},y_{t},z_{t})=$ $x_{t}^{\top}(\Theta_{t}^{Q}+Q)x_{t}+y_{t}^{\top}(\Theta_{t}^{u}+R^{u})y_{t}-z_{t}^{\top}(\Theta_{t}^{v}+R^{v})z_{t},$
$r^{2}(x_{t},y_{t},z_{t})+u^{2}_{t}(x_{t},y_{t},z_{t})=$ $x_{t}^{\top}(\Theta_{t}^{Q}-Q)x_{t}+y_{t}^{\top}(\Theta_{t}^{u}-R^{u})y_{t}-z_{t}^{\top}(\Theta_{t}^{v}-R^{v})z_{t}.$
Although the state, action and steering reward spaces are continuous, both the policies and steering reward are determined by their parameters. Therefore, the agents’ learning dynamics can be modeled by a function $f^{*}$ mapping between those parameters instead:
$(K_{t+1},L_{t+1})\sim f^{*}(\cdot|\underbrace{(K_{t},L_{t})}_{\text{agents' policies}},~{}\underbrace{(\Theta_{t}^{Q}+Q,\Theta_{t}^{u}+R^{u},\Theta_{t}^{u}+R^{v},\Xi_{t}^{Q}-Q,\Xi_{t}^{u}-R^{u},\Xi_{t}^{u}-R^{v})}_{\text{modified~{}rewards}}).$
Besides, the learning of steering strategy is equivalent to learning a function mapping $\psi$ from parameters in history $\{(K_{\tau},L_{\tau},\Theta_{\tau},\Xi_{\tau})\}_{\tau=1}^{t-1}\cup\{(K_{t},L_{t})\}$ to the next steering reward parameter $(\Theta_{t},\Xi_{t})$. Since both the policy parameters and steering reward parameters have finite dimension, this problem is tractable under our frameworks.
Uncountable function class $\mathcal{F}$
Our results can be extended to cases where the model class $\mathcal{F}$ is infinite but has a finite covering number. We denote $\mathcal{F}_{\varepsilon_{0}}$ as the $\varepsilon_{0}$-cover for $\mathcal{F}$, s.t.
$\forall f\in\mathcal{F},\quad\exists f^{\prime}\in\mathcal{F}_{\varepsilon_{0}},~{}\text{s.t.}~{}\max_{{\bm{\pi}}\in\Pi,{\bm{u}}}\mathbb{TV}\Big{(}f(\cdot|{\bm{\pi}},{\bm{u}})-f^{\prime}(\cdot|{\bm{\pi}},{\bm{u}})\Big{)}\leq\varepsilon_{0}.$
where $\mathbb{TV}$ denotes the total variation distance. If $\mathcal{F}$ is uncoutable but $\mathcal{F}_{\varepsilon_{0}}$ is finite, we run our algorithms with $\mathcal{F}_{\varepsilon_{0}}$ instead of $\mathcal{F}$. Under Assump. A, we denote $f^{*}_{\varepsilon_{0}}\in\mathcal{F}_{\varepsilon_{0}}$ is the function $\varepsilon_{0}$ close to $f^{*}$. By simmulation lemma, then we have:
$|\mathbb{E}_{\psi,f^{*}}[\eta^{\text{goal}}({\bm{\pi}}_{T+1})]-\mathbb{E}_{\psi,f^{*}_{\varepsilon_{0}}}[\eta^{\text{goal}}({\bm{\pi}}_{T+1})]|\leq$ $T\cdot\varepsilon_{0}\cdot\eta_{\max}$
$|\mathbb{E}_{\psi,f^{*}}[\sum_{t=1}^{T}\eta^{\text{cost}}({\bm{\pi}}_{t},{\bm{u}}_{t})]-\mathbb{E}_{\psi,f^{*}_{\varepsilon_{0}}}[\sum_{t=1}^{T}\eta^{\text{cost}}({\bm{\pi}}_{t},{\bm{u}}_{t})]|\leq$ $T^{2}\cdot\varepsilon_{0}\cdot\max_{{\bm{\pi}},{\bm{u}}}\eta^{\text{cost}}({\bm{\pi}},{\bm{u}}).$
As we can see, we can still optimize the objective in Eq. (1) with $\mathcal{F}_{\varepsilon_{0}}$, and then transfer guarantees on steering gap and cost for $f^{*}_{\varepsilon_{0}}$ (e.g. the the worst case guarantees in Prop. 3.3) to $f^{*}$ with additional $O(T^{2}\cdot\varepsilon_{0})$ errors, which is ignorable when $\varepsilon_{0}$ is small enough.
Non-Markovian Learning Dynamics
In general, non-Markovian learning dynamics is intractable, as implied by the fundamental difficulty in learning optimal policies in POMDPs. However, when some special structures exhibit, our methods for Markovian agents can be generalized. One example is the non-Markovian agents with finite-memory, i.e.,
${\bm{\pi}}_{t+1}\sim f^{*}(\cdot|{\bm{\pi}}_{t-m+1},{\bm{r}}+{\bm{u}}_{t-m+1},...,{\bm{\pi}}_{t},{\bm{r}}+{\bm{u}}_{t}).$
This can be reformulated by a Markovian dynamics
$\bm{x}_{t+1}\sim F^{*}(\cdot|{\bm{x}}_{t},{\bm{r}}+{\bm{u}}_{t}),$
with the same steering rewards as actions but a new definition of “state”: $\bm{x}_{t}:=\{{\bm{\pi}}_{t-m+1},{\bm{r}}+{\bm{u}}_{t-m+1},...,{\bm{\pi}}_{t-1},{\bm{r}}+{\bm{u}}_{t-1},{\bm{\pi}}_{t}\}$. Comparing with Def. 3.1, the dimension of the state space is expanded by $m$ times, which is still tractable for small $m$.
Neural Networks as Model Class to Approximate Complex $f^{*}$
The main principle for choosing $\mathcal{F}$ is to ensure our “realizability” assumption holds with high probability, i.e. the true model $f^{*}\in\mathcal{F}$. The concrete choice of $\mathcal{F}$ depends on the prior knowledge we have about the agents’ learning dynamics. In general, the less prior knowledge we have, the larger $\mathcal{F}$ should be to ensure realizability, and vice versa.
In practice, one “safe-choice” can be consider a class of parameterized neural networks as $\mathcal{F}$, since it has been proven in deep RL and supervised learning literature that neural networks have powerful approximation ability when $f^{*}$ is potentially very complex. Because in our formulation, we allow the randomness of next policy ${\bm{\pi}}_{t+1}$ (instead of a deterministic output) given ${\bm{\pi}}_{t}$ and ${\bm{r}}+{\bm{u}}_{t}$, we may consider a neural network taking the concatenation of ${\bm{\pi}}_{t}$, ${\bm{r}}+{\bm{u}}_{t}$ and another random Gaussian vector $\xi$ as inputs. Here the noise vector is introduced to model the stochasticity of ${\bm{\pi}}_{t}$.
The parameters of neural networks are in general continous variables, which implies the model class is uncountable. However, if the parameters has bounded value range, we can show the finite covering number on the parameter space. If we consider Lipschitz continuous activation functions (which is most of the cases), it implies the bounded covering number.
Besides, when considering neural networks, the resulting model class $\mathcal{F}$ can be extremely large and the MLE-based strategic exploration in Procedure 2 will be inefficient. We highlight that we design such exploration step in order to align with the main principle: the algorithm design should be supported by theoretical guarantees on the performance of the learned steering strategy. This focus on theoretical rigor is the main factor limiting the scalability of our algorithms in more complex settings. Conversely, if we relax the requirements on theoretical guarantees, it is not very challenging to adapt our algorithms to complex scenarios. For example, we can instead consider more scalable exploration methods, such as Random Network Distillation (RND) (Burda et al.,, 2018) or Bootstrapped DQN (Osband et al.,, 2016), although without theoretical guarantees.